OpenStudy (anonymous):

Evaluate: 32 + (6 − 2) • 4 − 6 over 3

OpenStudy (anonymous):

its out of 20, 23, 50, and 26

OpenStudy (anonymous):

Its 50

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

Yes because First you deal with the parentheses 32 + (7 - 2) x 4 - 6/3 = 32 + 5 x 4 - 6/3 Next you deal with the division and multiplication 32 + 5 x 4 - 6/3 = 32 + 20 - 2 Finally, you deal with the addition and subtraction in order 32 + 20 - 2 = 52 - 2 = 50

OpenStudy (anonymous):

okay thanks hun:)

OpenStudy (anonymous):

You're welcome